first five brothers were given land with area l03,107,105,90 and 95m². if the mean of the areas of land is 105m² who received the biggest share

1
Expert's answer
2021-11-17T12:45:51-0500

first brother's share: 107 m²

second brother's share: 105 m²

third brother's share: 90 m²

fourth brother's share: 95 m²

fifth brother's share: x


mean of the areas of land = 105 m²


"\dfrac{107 + 105 + 90 + 95 + x}5 = 105"


"397+x = 525\\
x = 525- 397\\
x = 128"


the fifth brother's share is therefore 128 m².


"\therefore" The fifth brother gets the largest share"."